The distance of a planet from a star is 10 AU. Express this distance in light years and parsec .
1
Expert's answer
2020-02-26T10:24:39-0500
"10\ AU=10(4.85\cdot 10^{−6})=4.85\cdot 10^{−5}\ parsec"

"10\ AU=10(15.8\cdot 10^{−6})=15.8\cdot 10^{−5}\ light-years"